OpenStudy (anonymous):

A ball rolls off the edge of a horizontal roof at 10 m/s. 2 seconds later, the speed of the ball will be? a. 10 m/s b. 19.6m/s c. 29.6 m/s d.22 m/s

OpenStudy (anonymous):

Assuming acceleration due to the gravity = 9.81m/s/s. Keep in mind that if the roof is horizontal, there is no initial vertical component to the ball's velocity - so u = 0. The ball could roll horizontally at 1000m/s and the initial vertical velocity would still be 0. Using the classical equations of motion (given the question, I'm assuming this is roughly where you're at in your physics studies?) v = u + at = 0 + 9.81*2 = 19.6m/s.